Q. 3) Carrying capacity / Stocking rate (Populations and ecosystem)
You removed the above-ground plant material from around 10 quadrats ( 1 sq meter each) in each area to measure primary production. After drying in the oven at 60° C for 24 hrs, the mean dry weight of the plant biomass was 290 g/m² in area 1 and 280 g/m² in area 2. If you want to keep a herd of wild onagers in a fenced area of 1 km², without feeding them during the spring season, what would be the maximum size of the herd (carrying capacity) in each area? You also estimated that around 30% of the plant mass in area 1 belong to unpalatable and poisonous plants, whereas the proportion of unpalatable plant material in area 2 was 25%.
Calculation hints (calculate the total amount of palatable plant biomass in 1 km² in each area, what is the average weight of an onager, how much dry mass does a herbivore mammal consume per day, …)
1
Expert's answer
2020-04-29T17:43:07-0400

"1)290\frac{g}{m^2};30 \%;\\m=\frac{290\times (100\%-30)}{100\%}=0.203kg;\\M=0.203\times1000000=203000kg\\
2)280\frac{g}{m^2};25 \%;\\m=\frac{280\times (100\%-25)}{100\%}=0.210kg;\\M=0.210\times1000000=210000kg"

  the average weight of an onager is 230kg - 6kg per day

"90d\times6=540kg;\\1)\frac{203000}{540}=375;\\2)\frac{210000}{540}=388;\\"

Answer:the first area - 375 onagers; the second area - 388 onagres